consider a slab with L=25cm A=90cm2 and the metal is copper .TH=125oc and Tc=10oc and steady state is reached,find the conduction rate through the slab.


1
Expert's answer
2022-06-06T14:48:58-0400

The Fourier's law says

"P=\kappa\frac{\Delta T}{\Delta h}A"

"P=400\:\rm \frac{W}{m\cdot K}*\frac{(125-10)\: K}{0.25\: m}*90*10^{-4}\: m^2"

"P=1656\:\rm W"